# Advancing the hBN Defects Database through Photophysical Characterization of Bulk hBN

### v1

@@ 2.3.2 Defect formation energy
where is the total energy of the supercell containing the defect in charge state ; is the total energy of pristine hBN; The term represents the number of atoms added or removed from the system, with being a chemical potential of atom species ; The Fermi level corresponds to the electronic chemical potential, referenced to the valence band maximum (VBM) of pristine hBN, denoted by . The final term, , is a correction for spurious electrostatic interactions arising from periodic boundary conditions in charged supercells [51]. In this work, we employed the Spinney Python code for charge correction [52].
@@ 2.4.5 Absorption and photoluminescence spectra
where represents the phonon-resolved Huang–Rhys factor as described in the previous section.
@@ 2.5 Database repository
GET https://raw.githubusercontent.com/ QCS-Theory/hBN-database/main/ hbn_defects_structure.db
@@ 3.3 Distribution of hBN defects
Lastly, turning to the radiative lifetime distribution shown in Fig. 4(c), the result suggests that most hBN defects exhibit lifetimes ranging from approximately 1 ns to 100 s. Furthermore, there is no clear correlation between the lifetime and the number of vacancies. More broadly, no systematic relationship is observed between the number of vacancies and other key defect properties, as further illustrated in Supplementary S1.
@@ 4 Programmatic Retrieval of the hBN Defect Database via the GitHub API
In your Python script or interactive session, import the functionfrom get_hBN_defects_database import get_database
@@
Invoke get_database() with the desired filtering criteriadata = get_database( option=["ZPL"], host=["monolayer", "bulk"], spin_multiplicity=["singlet", "doublet", "triplet"], charge_state=[-2, -1, 0, 1, 2], optical_spin_transition=["up", "down"], value_range=(2.0, 4.0), download_db=False )
The keyword arguments perform the following functions:
@@
The complete code examples are available in Supplementary S2.
@@ Acknowledgements.
- Lei et al. [2023] Y. Lei, Q. Lin, S. Xiao, J. Li, and H. Fang, Optically active telecom defects in mote2 fewlayers at room temperature, Nanomaterials 13, 10.3390/nano13091501 (2023).
@@
- Alkauskas et al. [2014] A. Alkauskas, Q. Yan, and C. G. Van de Walle, First-principles theory of nonradiative carrier capture via multiphonon emission, Phys. Rev. B 90, 075202 (2014).

### v2

@@ 2.3.2 Defect formation energy
where is the total energy of the supercell containing the defect in charge state ; is the total energy of pristine hBN; The term represents the number of atoms added or removed from the system, with being a chemical potential of atom species ; The Fermi level corresponds to the electronic chemical potential, referenced to the valence band maximum (VBM) of pristine hBN, denoted by . The final term, , is a correction for spurious electrostatic interactions arising from periodic boundary conditions in charged supercells [51]. In this work, we employed the Spinney Python code for charge correction [52]. Although charge states greater than 3 or less than 3 may be possible for certain defects, they were not considered in this work. Future studies may extend the analysis to include these regimes in order to identify additional stable defects and potential quantum emitters.
@@ 2.4.5 Absorption and photoluminescence spectra
where represents the phonon-resolved Huang–Rhys factor as described in the previous section. The convergence test of the supercell size is shown in Supplementary S1.
@@ 2.5 Database repository
GET https://raw.githubusercontent.com/
QCS-Theory/hBN-database/main/
hbn_defects_structure.db
@@ 3.3 Distribution of hBN defects
Lastly, turning to the radiative lifetime distribution shown in Fig. 4(c), the result suggests that most hBN defects exhibit lifetimes ranging from approximately 1 ns to 100 s. Furthermore, there is no clear correlation between the lifetime and the number of vacancies. More broadly, no systematic relationship is observed between the number of vacancies and other key defect properties, as further illustrated in Supplementary S2. Full example calculations of defect properties in the database are given in Supplementary S3.
@@ 4 Programmatic Retrieval of the hBN Defect Database via the GitHub API
In your Python script or interactive session, import the functionfrom get_hBN_defects_database import get_database -
@@
Invoke get_database() with the desired filtering criteriadata = get_database( option=["ZPL"], host=["monolayer", "bulk"], spin_multiplicity=["singlet", "doublet", "triplet"], charge_state=[-2, -1, 0, 1, 2], optical_spin_transition=["up", "down"], value_range=(2.0, 4.0), download_db=False )The keyword arguments perform the following functions:
@@
The complete code examples are available in Supplementary S4.
@@ 6 Conclusion
We stress that the primary goal of our work is to construct a large database covering comprehensive properties of many different defects. This is to serve as a screening tool that helps rule out unlikely defects and narrow down promising candidates. This work does not take into account symmetry reduction effects such as the Jahn–Teller distortion; therefore, the specific values of the ZPL may be shifted. Such effects should be carefully considered when determining a defect as a promising candidate. For instance, to use our database, if an experiment reports a ZPL at 2.2 eV, one should focus on defects with calculated ZPLs between 2.0 and 2.4 eV, taking into account the finite accuracy of DFT calculations. This approach effectively eliminates many candidates and suggests the most plausible ones.
@@ Acknowledgements.
- Lei et al. [2023] Y. Lei, Q. Lin, S. Xiao, J. Li, and H. Fang, Optically Active Telecom Defects in MoTe2 Fewlayers at Room Temperature, Nanomaterials 13, 1501 (2023).
@@
- Alkauskas et al. [2014] A. Alkauskas, B. B. Buckley, D. D. Awschalom, and C. G. Van De Walle, First-principles theory of the luminescence lineshape for the triplet transition in diamond NV centres, New J. Phys. 16, 073026 (2014).